# The Philosophical Foundations of Growing AI Like A Child

### v1

@@ 6 Conclusion
This paper proposes that the robustness challenge and Moravec’s Paradox, two key limitations of current LLMs with significant scientific and practical implications, may be jointly explained by differences in cognitive development between humans and machines. Specifically, these differences are not found within the dynamic process of improving representational power—a mechanism likely shared between scaling up in LLMs and growing up in humans. Instead, they stem from LLMs’ absence of core knowledge, a set of foundational cognitive abilities present in humans from early childhood. This core knowledge serves as the basis for gradually acquiring more complex skills over time. Empirical evidence presented in this paper demonstrates that such abilities are indeed missing in LLMs. Further analysis of the underlying causes of this core knowledge deficit suggests a viable solution: systematically increasing low-level representations of core cognitive domains within LLM training datasets.
This analysis underscores the need for future research to explore how core knowledge can be effectively incorporated into LLMs. Rather than contradicting the general principle of scaling laws, this perspective challenges the assumption that intelligence can emerge solely from domain-general cognitive mechanisms. Just as humans rely on developmental start-up software, LLMs may require structured early training to scaffold their cognitive growth. Encouragingly, given their ability to process vast amounts of high-level information, LLMs may not require innate structures but rather training data that prioritizes salient representations of low-level concepts—analogous to the perceptual input available to a child. Based on said theorization, this paper proposes an engineering solution leveraging large-scale synthetic data generation via simulation engines to systematically generate task scenarios based on developmental psychology paradigms. Importantly, there appears to be no fundamental technical barrier to pretraining core knowledge. The next step is to implement this approach and rigorously evaluate whether it enhances human-like cognitive competence, particularly in real-world robustness. This research agenda can be best summarized as growing AI like a child—at scale  [^4] The authors of this article are leading an engineering project on the conviction of this statement, and see https://growing-ai-like-a-child.github.io/ for details..

### v2

@@ 6 Conclusion
This paper proposes that the robustness challenge and Moravec’s Paradox, two key limitations of current LLMs with significant scientific and practical implications, may be jointly explained by differences in cognitive development between humans and machines. Specifically, these differences are not found within the dynamic process of improving representational power—a mechanism likely shared between scaling up in LLMs and growing up in humans. Instead, they stem from LLMs’ absence of core knowledge, a set of foundational cognitive abilities present in humans from early childhood. This core knowledge serves as the basis for gradually acquiring more complex skills over time. Empirical evidence presented in this paper demonstrates that such abilities are indeed missing in current models. Further analysis of the underlying causes of this core knowledge deficit suggests a viable solution: systematically diffusing low-level representations of core cognitive domains during LLM pretraining.
This analysis underscores the need for future research to explore how core knowledge can be effectively incorporated into LLMs. Rather than contradicting the general principle of scaling laws, this perspective challenges the assumption that intelligence can emerge solely from domain-general mechanisms. Just as humans rely on developmental start-up software, LLMs may require structured early training to scaffold their cognitive growth. Encouragingly, given their ability to process vast amounts of high-level information, LLMs may not require innate structures but rather training data that prioritizes salient representations of low-level concepts. Based on said theorization, this paper proposes an engineering solution leveraging large-scale synthetic data generation via simulation engines to systematically generate task scenarios based on developmental psychology paradigms. Importantly, there appears to be no fundamental technical barrier to pretraining core knowledge. The next step is to implement this approach and rigorously evaluate whether it enhances human-like cognitive competence, particularly in real-world robustness. This research agenda can be best summarized as growing AI like a child—at scale  [^5] The authors of this article are leading an engineering project on the conviction of this statement, and see http://grow-ai-like-a-child.com/ for details..